Steve fixed the PRM oplev pointing. I turned on the loops and measured the OLG, then set the pitch and yaw gains such that the upper UGF was ~8Hz (motivated by Jenne's loop design in ELOG 9401)
I then measured the oplev spectra of the optics as they were aligned for PRMI. (OSEMs on, oplevs on, LSC off, and ASC off)
Next, Jenne and I need to fix the ASC loop such that it properly accounts for the oplev loop.